# Union of Finitely Generated Congruences on Ground Term Algebra

### v1

@@ 14.1 Partial Depth First Search on the Auxiliary dpwpa
Observe that \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$ and \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$. As a result
@@
(\$$\left(\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}\;\mbox{or}\;\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}\right)=\mathrm{true}$.
@@
Algorithm PDFS examines the vertices of $AUX[E;F]$, it looks for vertices from which it can start the search. To this end, Algorithm PDFS decides whether for each $a\in C\langle E\cup F,\emptyset\rangle$, it holds that $a\in C\langle E,F\rangle$ or $a\in C\langle F,E\rangle$, that is, $\left(a.equal_{E}=\mathrm{true}\mbox{ or }a.equal_{F}=\mathrm{true}\right)$. First, Algorithm PDFS examines the vertex #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. It observes that #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$ and #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$. Then Algorithm PDFS starts the search from #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. It goes along the cycle (#$(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})$. Algorithm PDFS observes that #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$. Second, Algorithm PDFS examines the vertex \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$. Since \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, Algorithm PDFS does not start the search starting from \$$\$/_{\Theta\langle E\cup F,\emptyset\rangle}$. Consequently, Algorithm PDFS outputs true.
@@ 14.2 Decision Algorithm and its verification
    for each $a\in C\langle E\cup F,\emptyset\rangle$,
    • $a\in C\langle E,F\rangle$ or $a\in C\langle F,E\rangle$ and
    • if $a\not\in C\langle E,F\rangle$, then for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle F,E\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $b$.
@@
    for each $a\in C\langle E\cup F,\emptyset\rangle$,
    • $a.equal_{E}=\mathrm{true}$ or $a.equal_{F}=\mathrm{true}$ and
    • if $a\not\in C\langle E,F\rangle$, then for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{F}=\mathrm{true}$.
@@
    For each vertex $a\in C\langle E\cup F,\emptyset\rangle$, $a.equal_{E}=\mathrm{true}$ or $a.equal_{F}=\mathrm{true}$, and
@@
    Algorithm PDFS on $AUX[E;F]$ visits exactly once each vertex $a\in C\langle E\cup F,\emptyset\rangle$ such that there is a path of positive length or a cycle from some vertex $b\in C\langle E\cup F,\emptyset\rangle\setminus C\langle E,F\rangle$ to $a$. For each visited vertex $a$, $a.keeps_{F}=\mathrm{true}$.
@@
    There exists a vertex $a\in C\langle E\cup F,\emptyset\rangle$ such that $a.equal_{E}=\mathrm{true}$ or $a.equal_{F}=\mathrm{true}$.
@@
    Algorithm PDFS on $AUX[E;F]$ visits at most once each vertex $a\in C\langle E\cup F,\emptyset\rangle$ such that there is a path of positive length or a cycle from some vertex $b\in C\langle E\cup F,\emptyset\rangle\setminus C\langle E,F\rangle$ to $a$. Moreover, Algorithm PDFS visits a vertex $a\in C\langle E\cup F,\emptyset\rangle$ such that $a.keeps_{F}=\mathrm{false}$.
Proof.
By direct inspection of Definition 14.122 and by the equivalence $\mathrm{(iv)}\Leftrightarrow\mathrm{(vi)}$ in Lemma 14.127. ∎
@@
Proof.
By the equivalence $\mathrm{(i)}\Leftrightarrow\mathrm{(iv)}$ in Lemma 14.127 and by Lemma 14.128. ∎
@@
Recall that we have run Algorithm CAD on $\Sigma$, $E$, and $F$. By Proposition 4.47, we got the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$, and for every vertex $a\in C\langle E\cup F,\emptyset\rangle$, the values of the attributes $a.equal_{E}$, $a.equal_{F}$, $a.keeps_{E}$, and $a.keeps_{F}$. Then we run Algorithm PDFS on $AUX[E;F]$ in $O(n)$ time. Algorithm PDFS on input $AUX[E;F]$ outputs the answer to the question whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. Thence the complexity of our entire computation is $O(n\,\mathrm{log}\,n)$. ∎
@@
We now continue Examples 3.21, 7.73, 10.100, and 14.125. By the last lines of Example 10.100, GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together do not simulate the GTRS $R\langle E\cup F,\emptyset\rangle$. Accordingly Condition (iii) in Lemma 14.127 does not hold. For this reason, by Lemma 14.127, $\mathop{\leftrightarrow}^{*}_{E\cup F}\not\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. By the last line of Example, 14.125 Algorithm PDFS outputs false to the question whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. As a consequence, our decision algorithm also outputs false to the question whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
@@
We now continue Examples 3.22, 5.55, 7.74, 10.101, and 14.126. By the last line of Example 10.101, GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$. Therefore Condition (iii) in Lemma 14.127 holds. Consequently, by Lemma 14.127, $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. By the last line of Example, 14.126 Algorithm PDFS outputs true to the question whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. As a consequence, our decision algorithm also outputs true to the question whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
@@ 15 Main Case 4
In this main case, $E$ and $F$ are GTESs over a signature $\Sigma$ such that $\Sigma$ has a symbol of arity at least $2$, and GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are not total. We assume that we have already run Algorithm CAD on $\Sigma$, $E$, and $F$, and have obtained all the output it produces, see Proposition 4.47. Example 3.23 and its sequel Example 5.56, furthermore, Example 3.24 and its sequels Examples 5.57 and 15.134 demonstrate this case.
@@
We continue Examples 3.24 and 5.57. We recall that
$f(\#,\mathsterling)\mathop{\leftrightarrow}^{*}_{E\cup F}f(\$,\flat)$,  $(f(\#,\mathsterling),f(\$,\flat))\not\in\mathop{\leftrightarrow}^{*}_{E}$,   and   $(f(\#,\mathsterling),f(\$,\flat))\not\in\mathop{\leftrightarrow}^{*}_{F}$.
The idea of the ()($\Rightarrow$)( ⇒ ) direction of the proof of the next theorem comes from the proof of the above observation.
@@
()($\Rightarrow$)( ⇒ ) We proceed by contrapositive. Assume that $E\not\subseteq\mathop{\leftrightarrow}^{*}_{F}$ and $F\not\subseteq\mathop{\leftrightarrow}^{*}_{E}$. That is, there exists $l_{1}\doteq r_{1}\in E$ such that $(l_{1},r_{1})\not\in\mathop{\leftrightarrow}^{*}_{F}$ and there exists $l_{2}\doteq r_{2}\in F$ such that $(l_{2},r_{2})\not\in\mathop{\leftrightarrow}^{*}_{E}$. Then $l_{1}\neq r_{1}$ and $l_{2}\neq r_{2}$. Let $\sigma\in\Sigma_{k}$ with $k\geq 2$, and by Lemma 4.42, let $v\in T_{\Sigma}\setminus\bigcup W/_{\mathop{\leftrightarrow}^{*}_{E}}$, and $z\in T_{\Sigma}\setminus\bigcup W/_{\mathop{\leftrightarrow}^{*}_{F}}$. Then
@@
[…]
@@
[…]
@@
[…]
@@
()($\Leftarrow$)( ⇐ ) If $E\subseteq\mathop{\leftrightarrow}^{*}_{F}$, then $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{F}$, and hence $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. If $F\subseteq\mathop{\leftrightarrow}^{*}_{E}$, then $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}$, and hence $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. ∎
Consequence 15.137.
@@
By Consequence 15.137, we proceed as follows. Recall that we have run Algorithm CAD on $\Sigma$, $E$, and $F$. By Proposition 4.47, we got the relations $\rho\langle E,F\rangle$, $\rho\langle F,E\rangle$, and $\rho\langle E\cup F,\emptyset\rangle$ in $O(n\,\mathrm{log}\,n)$ time.
First we decide whether $E\subseteq\Theta\langle F,E\rangle$. By Consequence 15.137 and Proposition 4.25,
@@
We call above the functions $\mathrm{FIND}_{\rho\langle F,E\rangle}$ and $\mathrm{FIND}_{\rho\langle E,F\rangle}$ a total of at most $2n$ times, where $n=size(E)+size(F)$. These callings, by the main result of [16], require $O(n)$ time. As a result, the time complexity of our entire computation is $O(n\,\mathrm{log}\,n)$. ∎

### v2

@@ 14.1 Partial Depth First Search on the Auxiliary dpwpa
Observe that $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$ and $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$. As a result
@@
$\left(\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}\;\mbox{or}\;\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}\right)=\mathrm{true}$.
@@
Algorithm PDFS examines the vertices of $AUX[E;F]$, it looks for vertices from which it can start the search. To this end, Algorithm PDFS decides whether for each ${b}\in C\langle E\cup F,\emptyset\rangle$, it holds that ${b}\in C\langle E,F\rangle$ or ${b}\in C\langle F,E\rangle$, that is, $\left({b}.equal_{E}=\mathrm{true}\mbox{ or }a.equal_{F}=\mathrm{true}\right)$. First, Algorithm PDFS examines the vertex $_{\Theta\langle E\cup F,\emptyset\rangle}$. It observes that $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$ and $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$. Then Algorithm PDFS starts the search from $_{\Theta\langle E\cup F,\emptyset\rangle}$. It goes along the cycle $(\#/_{\Theta\langle E\cup F,\emptyset\rangle},\#/_{\Theta\langle E\cup F,\emptyset\rangle})$. Algorithm PDFS observes that $_{\Theta\langle E\cup F,\emptyset\rangle}.keeps_{F}=\mathrm{true}$. Second, Algorithm PDFS examines the vertex $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$. Since $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$, Algorithm PDFS does not start the search starting from $\mathdollar/_{\Theta\langle E\cup F,\emptyset\rangle}$. Consequently, Algorithm PDFS outputs true.
@@ 14.2 Decision Algorithm and its verification
    for each ${b}\in C\langle E\cup F,\emptyset\rangle$,
    • (a)
      ${b}\in C\langle E,F\rangle$ or ${b}\in C\langle F,E\rangle$ and
    • (b)
      if ${b}\not\in C\langle E,F\rangle$, then for each ${c}\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from b{b} to c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle F,E\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing c{c}.
@@
    for each ${b}\in C\langle E\cup F,\emptyset\rangle$,
    • (a)
      ${b}.equal_{E}=\mathrm{true}$ or ${b}.equal_{F}=\mathrm{true}$ and
    • (b)
      if ${b}\not\in C\langle E,F\rangle$, then for each ${c}\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from b{b} to c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then ${c}.keeps_{F}=\mathrm{true}$.
@@
We get the following lemma by direct inspection of Definition 14.132 and by the equivalence $\mathrm{(iv)}\Leftrightarrow\mathrm{(vi)}$ in Lemma 14.137.
@@
    For each vertex ${b}\in C\langle E\cup F,\emptyset\rangle$, ${b}.equal_{E}=\mathrm{true}$ or ${b}.equal_{F}=\mathrm{true}$, and
@@
    Algorithm PDFS on $AUX[E;F]$ visits exactly once each vertex ${b}\in C\langle E\cup F,\emptyset\rangle$ such that there is a path of positive length or a cycle from some vertex ${c}\in C\langle E\cup F,\emptyset\rangle\setminus C\langle E,F\rangle$ to b{b}. For each visited vertex b{b}, ${b}.keeps_{F}=\mathrm{true}$.
@@
    There exists a vertex ${b}\in C\langle E\cup F,\emptyset\rangle$ such that ${b}.equal_{E}=\mathrm{true}$ or ${b}.equal_{F}=\mathrm{true}$.
@@
    Algorithm PDFS on $AUX[E;F]$ visits at most once each vertex ${b}\in C\langle E\cup F,\emptyset\rangle$ such that there is a path of positive length or a cycle from some vertex ${c}\in C\langle E\cup F,\emptyset\rangle\setminus C\langle E,F\rangle$ to b{b}. Moreover, Algorithm PDFS visits a vertex ${b}\in C\langle E\cup F,\emptyset\rangle$ such that ${b}.keeps_{F}=\mathrm{false}$.
We get the following result by the equivalence $\mathrm{(i)}\Leftrightarrow\mathrm{(iv)}$ in Lemma 14.137 and by Lemma 14.138.
@@
Recall that we have run Algorithm CAD on $\Sigma$, $E$, and $F$. By Proposition 4.48, we got the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$, and for every vertex ${b}\in C\langle E\cup F,\emptyset\rangle$, the values of the attributes ${b}.equal_{E}$, ${b}.equal_{F}$, ${b}.keeps_{E}$, and ${b}.keeps_{F}$. Then we run Algorithm PDFS on $AUX[E;F]$ in $O(n)$ time. Algorithm PDFS on input $AUX[E;F]$ outputs the answer to the question whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. Thence the complexity of our entire computation is $O(n\,\mathrm{log}\,n)$. ∎
@@
We now continue Examples 3.22, 7.74, 10.110, and 14.135. By the last lines of Example 10.110, GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together do not simulate the GTRS $R\langle E\cup F,\emptyset\rangle$. Accordingly Condition (iii) in Lemma 14.137 does not hold. For this reason, by Lemma 14.137, $\mathop{\leftrightarrow}^{*}_{E\cup F}\not\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. By the last line of Example, 14.135 Algorithm PDFS outputs false to the question whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. As a consequence, our decision algorithm also outputs false to the question whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
@@
We now continue Examples 3.23, 5.56, 7.75, 10.111, and 14.136. By the last line of Example 10.111, GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$. Therefore Condition (iii) in Lemma 14.137 holds. Consequently, by Lemma 14.137, $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. By the last line of Example, 14.136 Algorithm PDFS outputs true to the question whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. As a consequence, our decision algorithm also outputs true to the question whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
@@ 15 Main Case 4
In this main case, $E$ and $F$ are GTESs over a signature $\Sigma$ such that $\Sigma$ has a symbol of arity at least $2$, and GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are not total. We assume that we have already run Algorithm CAD on $\Sigma$, $E$, and $F$, and have obtained all the output it produces, see Proposition 4.48. We show that $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if $E\subseteq\Theta\langle F,E\rangle$ or $F\subseteq\Theta\langle E,F\rangle$. By direct inspection of the relations $\tau\langle E,F\rangle$ and $\rho\langle F,E\rangle$, we decide whether $E\subseteq\Theta\langle F,E\rangle$. Similarly, by direct inspection of the relations $\tau\langle F,E\rangle$ and $\rho\langle E,F\rangle$, we decide whether $E\subseteq\Theta\langle F,E\rangle$. Example 3.24 and its sequel Example 5.57, furthermore, Example 3.25 and its sequels Examples 5.58 and 15.144 demonstrate this case.
@@
We continue Examples 3.25 and 5.58. We recall that
[…]
The idea of the $(\Rightarrow)$ direction of the proof of the next theorem comes from the proof of the above observation.
@@
$(\Rightarrow)$ We proceed by contrapositive. Assume that $E\not\subseteq\mathop{\leftrightarrow}^{*}_{F}$ and $F\not\subseteq\mathop{\leftrightarrow}^{*}_{E}$. That is, there exists $l_{1}\doteq r_{1}\in E$ such that $(l_{1},r_{1})\not\in\mathop{\leftrightarrow}^{*}_{F}$ and there exists $l_{2}\doteq r_{2}\in F$ such that $(l_{2},r_{2})\not\in\mathop{\leftrightarrow}^{*}_{E}$. Then $l_{1}\neq r_{1}$ and $l_{2}\neq r_{2}$. Let $\sigma\in\Sigma_{k}$ with $k\geq 2$, and by Lemma 4.43, let $v\in T_{\Sigma}\setminus\bigcup W/_{\mathop{\leftrightarrow}^{*}_{E}}$, and $z\in T_{\Sigma}\setminus\bigcup W/_{\mathop{\leftrightarrow}^{*}_{F}}$. Then
@@
be a sequence of $E$-rewrite steps. Then for each $i\in\{\,1,\ldots,n\,\}$, $\alpha_{i}$ is not a prefix of $11$.
@@
By induction on $n$, we show that for each $i\in\{\,1,\ldots,n\,\}$, $\alpha_{i}$ is not a prefix of $11$.
@@
Induction Step: Let $n\geq 2$ and let us assume that we have shown the statement for n1n-1. Then by the induction hypothesis, for each $i\in\{\,1,\ldots,n-1\,\}$, $\alpha_{i}$ is not a prefix of $11$. Accordingly we have $v=s_{0}|_{11}\mathop{\leftrightarrow}^{*}_{E}s_{n}|_{11}$. Consequently, by $v\not\in\bigcup W/_{\mathop{\leftrightarrow}^{*}_{E}}$, we have $s_{n}|_{11}\not\in\bigcup W/_{\mathop{\leftrightarrow}^{*}_{E}}$. Thus $s_{n}|_{11}$ is not a subterm of $p_{n}\in W$. As a consequence, $\alpha_{n}$ is not a prefix of $11$. Then the induction step is complete. ∎
Consider the sequence (3) of $E$ rewrite steps. By Claim 15.146, for each $i\in\{\,1,\ldots,n\,\}$, $\alpha_{i}$ is not a prefix of $11$. Consequently $t=\sigma(\sigma(t_{1},t_{2},\ldots t_{k}),p_{1},\ldots,p_{k-1})$ for some $p_{1},\ldots,p_{k-1}\in T_{\Sigma}$, and for each $j\in\{\,2,\ldots,k\,\}$, $l_{1}\mathop{\leftrightarrow}^{*}_{E}t_{j}$. Recall that $(l_{1},r_{1})\not\in\mathop{\leftrightarrow}^{*}_{E}$. Therefore for each $j\in\{\,2,\ldots,k\,\}$, $t_{j}\neq r_{1}$. Hence
@@
$(\Leftarrow)$ If $E\subseteq\mathop{\leftrightarrow}^{*}_{F}$, then $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{F}$, and hence $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. If $F\subseteq\mathop{\leftrightarrow}^{*}_{E}$, then $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}$, and hence $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. ∎
Consequence 15.147.
@@
By Consequence 15.147, we proceed as follows. Recall that we have run Algorithm CAD on $\Sigma$, $E$, and $F$. By Proposition 4.48, we got the relations $\rho\langle E,F\rangle$, $\rho\langle F,E\rangle$, and $\rho\langle E\cup F,\emptyset\rangle$ in $O(n\,\mathrm{log}\,n)$ time.
First we decide whether $E\subseteq\Theta\langle F,E\rangle$. By Proposition 4.26,
@@
By Consequence 15.147
We call above the functions $\mathrm{FIND}_{\rho\langle F,E\rangle}$ and $\mathrm{FIND}_{\rho\langle E,F\rangle}$ a total of at most $2n$ times, where $n=\mathrm{size}(E)+\mathrm{size}(F)$. These callings, by the main result of [3], require $O(n)$ time. As a result, the time complexity of our entire computation is $O(n\,\mathrm{log}\,n)$. ∎